Robert is now 1/2 his dad's age. Twelve years ago, robert was 1/3 his dad's age.
Find both their ages now? Can you please help. I messed 4 days of school because of the flu.

Answer by checkley77(12844)

Dad=x
Robert=x/2
x/2-12=x/3
x/2-x/3=12
(3x-2x)/6=12
x/6=12
x=72 age of Dad now.
72/2=36 Robert's age now.
Proof:
36-12=72/3
24=24